# Behavior-Regularized Diffusion Policy Optimization for Offline Reinforcement Learning

### v1

@@ 4.2 Actor-Critic across Two Time Scales
$\mathcal{B}^{\pi}Q^{\pi}(s,a)=R(s,a)+\gamma\mathbb{E}\Big{[}Q^{\pi}(s^{\prime},a^{0})-\eta\sum_{n=1}^{N}\ell^{\pi,s^{\prime}}_{n}(a^{n})\Big{]},$ (12)
i.e., we sample diffusion paths $a^{0:N}$ at the next state $s^{\prime}$, calculate the Q-values $Q^{\pi}(s^{\prime},a^{0})$ and the accumulated penalties along the path, and perform a one-step TD backup. For the diffusion policy, its objective can be expressed as:
@@
At first glance, one might treat the pathwise KL as a whole, and optimize the diffusion policy by back-propagating the gradient throughout the sampled path $a^{0:N}$, similar to Diffusion-QL (Wang et al., 2023). Nevertheless, this requires preserving the computation graph of all diffusion steps and therefore incurs considerable computational overhead. Moreover, the inherent stochasticity of the diffusion path introduces significant variance during optimization, which can hinder the overall performance.
@@
Intuitively, $V^{\pi,s}_{n}(a^{n})$ receives the state $s$, intermediate generation result $a^{n}$, and the diffusion step $n$ as input, and estimates the expected accumulative penalties of continuing the diffusion path at $a^{n}$ and step $n$. It operates completely inside each environment timestep and performs TD updates across the diffusion steps. The semantic illustration of the value function backups is presented in Figure 3.
@@
Input: Offline dataset $\mathcal{D}$, initialized diffusion policy $p^{\pi}$, Q-value networks $Q_{\psi_{k}}$, diffusion value networks $V_{\phi_{k}}$.
@@
2:  Pretrain the reference diffusion policy $p^{\nu}$ via Eq. (9)
@@
5:  for $i=1,2,\dots,N_{\text{total\_steps}}$ do
6:     Sample minibatch $B=\{(s_{t},a_{t},s^{\prime}_{t},r_{t})\}_{t=1}^{|B|}$ from $\mathcal{D}$
@@
8:     Extend each $(s_{t},a_{t},s^{\prime}_{t},r_{t})$ in $B$ by sampling $n\sim U[1,N]$ and $a^{n}_{t}\sim q_{n|0}(\cdot|a_{t})$
@@
11:        Update $p^{\pi}$ by performing gradient ascend with the objective defined in Eq. (15)
@@
Let $\Pi$ be a set of admissible policies satisfying $\forall\,\pi\in\Pi,\;\forall\,s\in\mathcal{S}$,
$\sup_{a^{0}\in\mathcal{A}}\;\frac{\pi(a^{0}\mid s)}{\nu(a^{0}\mid s)}=\frac{\int p^{\pi,s}(a^{0:N})\mathrm{d}a^{1:N}}{\int p^{\nu,s}(a^{0:N})\mathrm{d}a^{1:N}}<\infty,$
where $p^{\nu}$ is a reference diffusion policy.
@@
(Policy Improvement) Let $p^{\pi_{\textrm{new}}}$ be the optimizer of the optimization problem defined in Eq. (15). Then, $V^{\pi_{\textrm{new}},s}_{n}(a)\geq V^{\pi_{\textrm{old}},s}_{n}(a)$ under Assumption 4.3.
Proof.
Proof can be found in Appendix C.4. ∎
Combining Proposition 4.4 and a similar convergence argument in Soft Actor Critic (Haarnoja et al., 2018), we obtain that the policy is guaranteed to improve w.r.t. the pathwise KL problem defined in 4.1.
@@
(Policy Iteration) Under Assumption 4.3, repeated application of soft policy evaluation in Eq. (12) and Eq. (14) and soft policy improvement in Eq. (15) from any $p^{\pi}\in\Pi$ converges to a policy $p^{\pi^{*}}$ such that $V_{n}^{\pi^{*},s}(a)\geq V_{n}^{\pi,s}(a)$ for all $p^{\pi}\in\Pi$, $n\in\{0,1,\ldots,N\}$, and $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Proof.
Proof can be found in Appendix C.5. ∎
Dataset CQL IQL DD SfBC IDQL-A QGPO DTQL Diffusion-QL DAC BDPO (Ours)
halfcheetah-m 44.0 47.4 49.1 45.9 51.0 54.1 57.9 51.1 59.1 69.9$\pm$0.4
hopper-m 58.5 66.3 79.3 57.1 65.4 98.0 99.6 90.5 101.2 99.3$\pm$2.9
walker2d-m 72.5 78.3 82.5 77.9 82.5 86.0 89.4 87.0 96.8 96.0$\pm$1.2
halfcheetah-m-r 45.5 44.2 39.3 37.1 45.9 47.6 50.9 47.8 55.0 53.9$\pm$1.8
hopper-m-r 95.0 94.7 100.0 86.2 92.1 96.9 100.0 101.3 103.1 102.2$\pm$0.4
walker2d-m-r 77.2 73.9 75.0 65.1 85.1 84.4 88.5 95.5 96.8 99.0$\pm$0.6
halfcheetah-m-e 91.6 86.7 90.6 92.6 95.9 93.5 92.7 96.8 99.1 104.7$\pm$3.9
hopper-m-e 105.4 91.5 111.8 108.6 108.6 108.0 109.3 111.1 111.7 111.4$\pm$0.1
walker2d-m-e 108.8 109.6 108.8 109.8 112.7 110.7 110.0 110.1 113.6 115.7$\pm$0.3
locomotion sum 698.5 749.7 736.2 680.3 739.2 779.4 798.3 791.2 836.4 852.1
antmaze-u 74.0 87.5 - 92.0 94.0 94.8 96.4 93.4 99.5 99.8$\pm$0.4
antmaze-u-div 84.0 62.2 - 85.3 80.2 74.4 78.8 66.2 85.0 80.2$\pm$6.4
antmaze-m-play 61.2 71.2 - 81.3 84.2 83.6 79.6 76.6 85.8 88.4$\pm$3.3
antmaze-m-div 53.7 70.0 - 82.0 84.8 83.8 82.2 78.6 84.0 85.0$\pm$5.9
antmaze-l-play 15.8 39.6 - 59.3 63.5 66.6 52.0 46.4 50.3 69.6$\pm$15.3
antmaze-l-div 14.9 47.5 - 45.5 67.9 64.8 66.4 56.6 55.3 79.0$\pm$3.0
antmaze sum 303.6 378.0 - 445.4 474.6 467.9 455.4 417.8 459.9 502.0
Remark (Actor-critic across two time scales). Our treatment of the environment MDP and the (reverse) diffusion MDP resembles DPPO (Ren et al., 2024), which unfolds the diffusion steps into the environment MDP and employs the PPO (Schulman et al., 2017) algorithm to optimize the diffusion policy in the extended MDP. Compared to DPPO, our method adds penalty in each diffusion step to implement behavior regularization. Besides, we maintain two types of value functions to account for both inter-timestep and inner-timestep policy evaluation, thereby offering low-variance policy gradients.
@@ 4.3 Practical Implementation
$\operatorname{KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]=\frac{\|\mu^{\pi,s}_{n}(a^{n})-\mu^{\nu,s}_{n}(a^{n})\|^{2}}{2\sigma_{n}^{2}}.$ (16)
That is, each penalty term is simply the discrepancy between the mean vectors of the reverse transitions, weighted by a factor depending on the noise schedule. The transition of the reverse process becomes Gaussian exactly in the continuous limit, in which case, the KL divergence can be computed through Girsanov’s theorem (Oksendal, 2013), and the sum of mean squared errors (MSE) is replaced by an integral (Franzese et al., 2024). See appendix D for more details.
Selection of States, Diffusion Steps and Actions. The policy improvement step requires training $V^{\pi,s}_{n}$ and improving $p^{\pi}$ on every $(s,a^{n},n)$ triplet sampled from the on-policy distribution of $p^{\pi}$. However, we employ an off-policy approach, by sampling $(s,a)$ from the dataset, $n$ according to the Uniform distribution $U[1,N]$ and $a^{n}$ according to $q_{n|0}(\cdot|a)$, which we found works sufficiently well in our experiments. We note that on-policy sampling $(s,a^{n},n)$ may further benefits the performance at the cost of sampling and restoring these triplets on the fly, similar to what iDEM (Akhound-Sadegh et al., 2024) did in its experiments.
Figure 4: Generation paths of BDPO on 8gaussian task (top) and 2spirals (down) task. The regularization strength is set to $\eta=0.06$, which is identical to Figure 5. The first four columns depict the diffusion generation process in different time intervals, with the green dots being the starting points of those intervals, the red dots being the ending points, and the grey lines in between are intermediate samples. The rightmost figures depict the final action samples. We use DDIM sampling for better illustration.
@@
where $\bar{\phi}_{k}$ and $\bar{\psi}_{k}$ are exponentially moving average versions of the value networks. The objective for value networks is minimizing the mean-squared error between the prediction and the target:
($\mathcal{L}(\{\psi_{k}\}_{k=1}^{K})$ $=\mathbb{E}_{(s,a,s^{\prime},r)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{Q}-Q_{\psi_{i}}(s,a))^{2}\right],$ (18)
($\mathcal{L}(\{\phi_{k}\}_{k=1}^{K})$ $=\mathbb{E}_{(s,a)\sim\mathcal{D},n,a^{n}}\left[\sum_{k=1}^{K}(y_{V}-V^{\pi,s}_{n,\phi_{i}}(s,a))^{2}\right],$
@@ 5 Experiments
Figure 5: Illustration of the 8gaussians and 2spirals dataset and the re-sampled result with $\eta=0.06$.
We evaluate BDPO with synthetic 2D tasks and also the D4RL benchmark (Fu et al., 2020). Due to the space limit, a detailed introduction about these tasks and the hyper-parameter configurations are deferred to Section A and Section B. More experimental results, such as the comparison of computational efficiency, are deferred to Appendix E.

### v2

@@ 4.2 Actor-Critic across Two Time Scales
[…]
i.e., we sample diffusion paths $a^{\prime 0:N}$ at the next state $s^{\prime}$, calculate the Q-values $Q^{\pi}(s^{\prime},a^{\prime 0})$ and the accumulated penalties along the path, and perform a one-step TD backup. For the diffusion policy, its objective can be expressed as:
@@
At first glance, one might treat the pathwise KL as a whole and optimize the diffusion policy by back-propagating the gradient throughout the sampled path $a^{0:N}$, similar to Diffusion-QL (Wang et al., 2023). Nevertheless, this requires preserving the computation graph of all diffusion steps and therefore incurs considerable computational overhead. Moreover, the inherent stochasticity of the diffusion path introduces significant variance during optimization, which can hinder the overall performance.
@@
Intuitively, $V^{\pi,s}_{n}(a^{n})$ receives the state $s$, intermediate generation result $a^{n}$, and the diffusion step $n$ as input, and estimates the expected cumulative penalties of continuing the diffusion path at $a^{n}$ and step $n$. It operates completely inside each environment timestep and performs TD updates across the diffusion steps. The semantic illustration of the value function backups is presented in Figure 3. The following proposition demonstrates the convergence of policy evaluation.
@@
Input: Offline dataset $\mathcal{D}$, initialized diffusion policy $p^{\pi}$ and $p^{\nu}$, Q-value networks $Q_{\psi_{k}}$, diffusion value networks $V_{\phi_{k}}$.
@@
2:  Pretrain the behavior diffusion policy $p^{\nu}$ via Eq. (9)
@@
5:  for $i=1,2,\cdots,N_{\text{total\_steps}}$ do
6:     Sample minibatch $B=\{(s_{i},a_{i},s^{\prime}_{i},r_{i})\}_{i=1}^{|B|}$ from $\mathcal{D}$
@@
8:     Extend each $(s_{i},a_{i},s^{\prime}_{i},r_{i})$ in $B$ by sampling $n\sim U[1,N]$ and $a^{n}_{t}\sim q_{n|0}(\cdot|a_{t})$
@@
11:        Update $p^{\pi}$ by performing gradient ascend with the objective defined in Eq. (16) and $B$
@@
Let $\Pi$ be the set of admissible diffusion policies satisfying $\forall\,p^{\pi}\in\Pi,\;\forall\,s\in\mathcal{S}$,
$\sup_{a^{0}\in\mathcal{A}}\;\frac{\pi(a^{0}|s)}{\nu(a^{0}|s)}=\frac{\int p_{0:N}^{\pi,s}(a^{0:N})\mathrm{d}a^{1:N}}{\int p_{0:N}^{\nu,s}(a^{0:N})\mathrm{d}a^{1:N}}<\infty,$
where $p^{\nu}$ is the behavior diffusion policy.
@@
(Soft Policy Improvement, proof in Appendix C.4) Let $p^{\pi_{\textrm{new}}}$ be the optimizer of the problem defined in Eq. (16). Under Assumption 4.3, $V^{\pi_{\textrm{new}},s}_{n}(a^{n})\geq V^{\pi_{\textrm{old}},s}_{n}(a^{n})$ holds for all $n\in\{0,1,\ldots,N\}$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Combining Proposition 4.4 and a similar convergence argument in Soft Actor Critic (Haarnoja et al., 2018), we obtain that the policy is guaranteed to improve w.r.t. the pathwise KL problem in Definition 4.1.
@@
(Policy Iteration, proof in Appendix C.5) Under Assumption 4.3, repeated application of soft policy evaluation in Eq. (13) and Eq. (15) and soft policy improvement in Eq. (16) from any $p^{\pi}\in\Pi$ converges to a policy $p^{\pi^{*}}$ such that $V_{n}^{\pi^{*},s}(a)\geq V_{n}^{\pi,s}(a)$ for all $p^{\pi}\in\Pi$, $n\in\{0,1,\ldots,N\}$, and $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Figure 4: Generation paths of BDPO on the 8gaussian (top), 2spirals (middle), and moons (down) datasets. The regularization strength is set to $\eta=0.06$, which is identical to Figure 5. The first five columns depict the diffusion generation process at different time intervals, with green dots indicating the starting points of these intervals, red dots indicating the ending points, and grey lines in between representing intermediate samples. The background color depicts the output from the diffusion value functions $V^{\pi,s}_{n}$ in the entire 2D space. The rightmost figures depict the final action samples. We use DDIM sampling (Song et al., 2020) for better illustration.
Figure 5: Illustration of the 8gaussians, 2spirals, and moons datasets. The top row depicts the original data distribution $p_{\text{data}}$, while the second row depicts the target distribution $p_{\text{target}}$ at $\eta=0.06$ by re-sampling data points according to their energies.
Remark (Actor-critic across two time scales). Our treatment of the environment MDP and the (reverse) diffusion MDP resembles DPPO (Ren et al., 2024), which unfolds the diffusion steps into the environment MDP and employs the PPO (Schulman et al., 2017) algorithm to optimize the diffusion policy in the extended MDP. Compared to DPPO, our method adds a penalty in each diffusion step to implement behavior regularization. Besides, we maintain two types of value functions to account for both inter-timestep and inner-timestep policy evaluation, thereby offering low-variance policy gradients.
@@ 4.3 Practical Implementation
$D_{\rm KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]=\frac{\|\mu^{\pi,s}_{n}(a^{n})-\mu^{\nu,s}_{n}(a^{n})\|^{2}}{2\sigma_{n}^{2}}.$ (17)
That is, each penalty term is simply the discrepancy between the mean vectors of the reverse transitions, weighted by a factor depending on the noise schedule. The transition of the reverse process becomes Gaussian exactly in the continuous limit, in which case, the KL divergence can be computed through Girsanov’s theorem (Oksendal, 2013), and the sum of mean squared errors (MSEs) is replaced by an integral (Franzese et al., 2024). In Appendix D, we demonstrate that the pathwise KL is consistent with the result from Girsanov’s theorem in the limit of infinitesimal diffusion step size.
Dataset CQL IQL DD SfBC IDQL-A QGPO SRPO DTQL Diffusion-QL EDP DAC BDPO (Ours)
halfcheetah-m 44.0 47.4 49.1 45.9 51.0 54.1 60.4 57.9 51.1 52.1 59.1 71.2$\pm$0.9
hopper-m 58.5 66.3 79.3 57.1 65.4 98.0 95.5 99.6 90.5 81.9 101.2 100.6$\pm$0.7
walker2d-m 72.5 78.3 82.5 77.9 82.5 86.0 84.4 89.4 87.0 86.9 96.8 93.4$\pm$0.5
halfcheetah-m-r 45.5 44.2 39.3 37.1 45.9 47.6 51.4 50.9 47.8 49.4 55.0 58.9$\pm$0.9
hopper-m-r 95.0 94.7 100.0 86.2 92.1 96.9 101.2 100.0 101.3 101.0 103.1 101.4$\pm$0.5
walker2d-m-r 77.2 73.9 75.0 65.1 85.1 84.4 84.6 88.5 95.5 94.9 96.8 95.5$\pm$1.6
halfcheetah-m-e 91.6 86.7 90.6 92.6 95.9 93.5 92.2 92.7 96.8 95.5 99.1 108.7$\pm$0.9
hopper-m-e 105.4 91.5 111.8 108.6 108.6 108.0 100.1 109.3 111.1 97.4 111.7 111.3$\pm$0.2
walker2d-m-e 108.8 109.6 108.8 109.8 112.7 110.7 114.0 110.0 110.1 110.2 113.6 115.6$\pm$0.4
locomotion sum 698.5 749.7 736.2 680.3 739.2 779.2 783.8 798.3 791.2 769.3 836.4 856.7
antmaze-u 74.0 87.5 - 92.0 94.0 96.4 90.8 94.8 93.4 94.2 99.5 98.4$\pm$1.5
antmaze-u-div 84.0 62.2 - 85.3 80.2 74.4 59.0 78.8 66.2 79.0 85.0 83.0$\pm$12.4
antmaze-m-play 61.2 71.2 - 81.3 84.5 83.6 73.0 79.6 76.6 81.8 85.8 92.0$\pm$2.9
antmaze-m-div 53.7 70.0 - 82.0 84.8 83.8 65.2 82.2 78.6 82.3 84.0 93.2$\pm$2.8
antmaze-l-play 15.8 39.6 - 59.3 63.5 66.6 38.8 52.0 46.4 42.3 50.3 69.0$\pm$8.0
antmaze-l-div 14.9 47.5 - 45.5 67.9 64.8 33.8 54.0 56.6 60.6 55.3 84.0$\pm$3.2
antmaze sum 303.6 378.0 - 445.4 474.9 469.6 360.6 441.4 417.8 440.2 459.9 519.6
Selection of States, Diffusion Steps and Actions. The policy improvement step requires training $V^{\pi,s}_{n}$ and improving $p^{\pi}$ on every $(s,a^{n},n)$ triplet sampled from the on-policy distribution of $p^{\pi}$. However, we employ an off-policy approach, by sampling $(s,a)$ from the dataset, $n$ according to the Uniform distribution $U[1,N]$ and $a^{n}$ according to $q_{n|0}(\cdot|a)$, which we found works sufficiently well in our experiments. We note that on-policy sampling $(s,a^{n},n)$ may further benefit the performance at the cost of sampling and restoring these triplets on the fly, similar to what iDEM (Akhound-Sadegh et al., 2024) did in its experiments.
@@
where $\bar{\phi}_{k}$ and $\bar{\psi}_{k}$ are exponentially moving average versions of the value networks, and $\mathrm{Avg},\mathrm{Var}$ denote averages and variances respectively. The objectives for value networks are to minimize the mean-squared error between the prediction and the target:
($\mathcal{L}(\{\psi_{k}\}_{k=1}^{K})$ $=\mathbb{E}_{(s,a,s^{\prime},r)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{Q}-Q^{\pi}_{\psi_{k}}(s,a))^{2}\right],$ (19)
($\mathcal{L}(\{\phi_{k}\}_{k=1}^{K})$ $=\mathbb{E}_{(s,a)\sim\mathcal{D},n,a^{n}}\left[\sum_{k=1}^{K}(y_{V}-V^{\pi,s}_{n,\phi_{k}}(a^{n}))^{2}\right],$
@@ 5 Experiments
We evaluate BDPO with synthetic 2D tasks and also the D4RL benchmark (Fu et al., 2020). Due to the space limit, a detailed introduction about these tasks and the hyperparameter configurations is deferred to Section A and Section B. More experimental results are deferred to Appendix E.
Figure 6: Sensitivity analysis of the regularization strength $\eta$. For each configuration, we report the mean and the standard deviation of the performances aggregated from 5 independent seeds and 10 evaluation episodes for each seed.
Figure 7: Sensitivity analysis of the lower confidence bound coefficient $\rho$. For each configuration, we report the mean and the standard deviation of the performances aggregated from 5 independent seeds and 10 evaluation episodes for each seed.
5.1 Results of Synthetic 2D Datasets
We leverage synthetic 2D datasets used in Lu et al. (2023) to further our understanding of the mechanism of BDPO. Each dataset contains data points $x_{i}$ paired with specific energy values $\mathcal{E}(x_{i})$. By re-sampling the data according to the Boltzmann distribution $p(x_{i})\propto\exp(\mathcal{E}(x_{i})/\eta)$, we can obtain the ground truth target distribution at a specific regularization strength $\eta$ (Haarnoja et al., 2018; Nair et al., 2020):
$p_{\text{target}}(x_{i})\propto p_{\text{data}}(x_{i})\exp(\mathcal{E}(x_{i})/\eta).$
As an example, Figure 5 illustrates both the original dataset and the re-sampled dataset with $\eta=0.06$ for the 8gaussians, 2spirals, and moons datasets.
Our objective is to analyze the properties of the diffusion policy and the diffusion value function trained with BDPO. The results are depicted in Figure 4 (full results on all datasets in Figure 13), where each row corresponds to a different dataset and visualizes the iterative sampling process. From left to right, the samples evolve from an initial noisy distribution toward a well-defined structure as the reverse diffusion progresses. In the initial steps ($n=50$ to $n=40$), the sample movement is subtle, while in the later steps ($n=30$ to $n=0$), the samples rapidly converge to the nearest modes of the data. The final column on the right presents the ultimate samples generated by the policy, which closely match the ground-truth target distribution visualized in Figure 5.
Besides the generation result, the background color in Figure 4 depicts the landscapes of the diffusion value functions $V^{\pi}$ at $n=50/40/30/20/10/0$. In earlier, noisier steps (e.g., $n=50$), the output values across the space are similar, resulting in a smoother landscape and weaker guidance during sampling. As the noise level decreases in later steps (e.g., $n=10$), the value outputs vary more sharply, creating stronger guidance on the diffusion policy. Such progression enables the model to refine samples while exploring sufficiently across the space, finally yielding results that align with the target distribution.
5.2 Results on Continuous Control Tasks
To assess the performance of BDPO on complex continuous control problems, we utilize datasets from D4RL as our test bench and compare BDPO against a wide spectrum of offline RL algorithms. Specifically, we include 1) CQL (Kumar et al., 2020b) and IQL (Kostrikov et al., 2022), which are representatives using straightforward one-step policies; 2) Decision Diffuser (DD) (Ajay et al., 2023), which uses diffusion models for planning and extracts actions from planned trajectories; 3) SfBC (Chen et al., 2023), IDQL (Hansen-Estruch et al., 2023), and QGPO (Lu et al., 2023), which use diffusion to model the behavior and further refine actions using the value functions; 4) SRPO (Chen et al., 2024a) and DTQL (Chen et al., 2024b), which use diffusion to provide behavior regularization for one-step actors; 5) Diffusion-QL (Wang et al., 2023), which propagates the gradient of Q-value functions over the whole generation process; and 6) EDP (Kang et al., 2023) and DAC (Fang et al., 2024), which accelerate the training of diffusion policies by using action approximation or using proxy objectives. The performance of each algorithm is measured by the normalized score on each task (see Appendix A for details).
The results are listed in Table 1. Our findings indicate that diffusion-based methods, particularly those with diffusion-based actor and regularization (including BDPO, DAC, and Diffusion-QL), substantially outperform their non-diffusion counterparts, especially in locomotion tasks. Meanwhile, BDPO consistently achieves superior performance across nearly all datasets, underscoring the effectiveness of combining diffusion policies and the behavior-regularized RL framework. Besides the final performance, we plot the training curves of BDPO in Section E.4. Overall, BDPO achieves fast and stable convergence except for some of the antmaze datasets, where variations may occur due to the sparse reward nature of these datasets.